OpenStudy (anonymous):

Will give medal! Find the third, sixth, and twentieth terms of the sequence described by the following rule. Be sure to show all your work. A(n) = 12 + (n – 1)(-4)

OpenStudy (anonymous):

The rule simplified to: 12 - 4n + 4 = 16 - 4n So 16-4n will give you any term. 3rd term, let n = 3, you get 16 - 4(3) = 16 - 12 = 4 6th term, let n = 6, you get 16 - 4(6) - 16 - 24 = -8 You can now compute the 20th term or any term you want.

OpenStudy (anonymous):

Simply use 3,6 and 20 as value of n. So A(3)=12+(3-1)*(-4)=12+2*(-4)=12-8=4 A(6)=12+(6-1)(-4)=12+5(-4)=12-20=-8 A(20)=12+(20-1)(-4)=12+19(-4)=12-76=-64